# A Concentration Bound for TD(0) with Function Approximation

### v2

@@ 1 Introduction
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2023), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$). Korda and Prashanth (2015) give a high probability bound for the same online framework as us.
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2023), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
@@ 2 TD(0) - Background and Main Result
TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies (Tsitsiklis and Van Roy 1997). Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $S$. The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed. Assume that the chain is irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$. Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$. Reward $r(i)$ is received when a transition from state $i$ takes place. The objective is to evaluate the long term discounted reward for each state given by the “value function”
@@
    and is assumed to be non-increasing, i.e., $a(n+1)\leq a(n)\ \forall\;n$. We also assume that $a(n)<1$ for all $n$. We further assume that $\frac{d_{1}}{n}\leq a(n)\leq d_{3}\left(\frac{1}{n}\right)^{d_{2}},\forall\ n$, where $d_{1}>0$ and $0<d_{2}\leq 1$. Larger values of $d_{1}$ and $d_{2}$ and smaller values of $d_{3}$ improve the main result presented below. The role this assumption plays in our bounds will become clear later.
@@
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor. Prior works such as Tsitsiklis and Van Roy (1997) show the contractive nature with respect to the weighted norm $\|\cdot\|_{D}$.
@@
The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$. This can be bounded using high probability bounds for TD(0) (e.g., Korda and Prashanth (2015)). Alternatively, moment bounds (Srikant and Ying 2019, Bhandari et al. 2018, Chen et al. 2021) can be combined with Markov’s inequality to bound this term.
@@
For the special case of $a(n)=\frac{d_{1}}{n}$, and by combining our result with the finite time bound in Prashanth et al. (2021), we can rewrite the result in Theorem 1 in the following form.
@@
Let $a(n)=d_{1}/n$ with sufficiently large $d_{1}$, and $n_{0}$ large enough to satisfy assumptions of Theorem 1. Then with probability at least $1-\varepsilon$, we have for all $m\geq n_{0}$ that
$\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
The proof for this corollary has been presented at the end of Appendix B. Chen et al. (2023) gives an all-time bound which matches our result at $n_{0}$, but their bound goes to zero as $m\uparrow\infty$. In our bound, the term $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ remains constant as $m$ is increased. Here $\delta$ can be modified to $\delta(m)$ (similar to the treatment in Corollary 1 in Chandak et al. (2022)). But the term $a(n_{0})(c_{2}+c_{1}\epsilon)$ arises from our treatment of Markov noise using Poisson equation. Note that Chen et al. (2023) do not consider the case of Markov noise, but only consider the case of additive and multiplicative noise (i.i.d. samples of state-action-next state triplets). We leave incorporating their ideas into our approach to get a bound decaying with $m$ for Markov noise as a future work.
@@ 3 Proof of the Main Result
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction. Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ for $n_{0}\leq m\leq n-1$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$. Then,
@@
Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1.
@@ 4 Conclusions
In conclusion, we note some future directions. The concept of relaxed martingale concentration inequalities to obtain bounds of the similar flavor for algorithms which suffer from similar issues. These include TD($\lambda$) and SSP Q Learning. Alternatively, similar bounds can be obtained for variants of temporal difference learning (Chen et al. 2021).
@@ B.4 Proof of Corollary 1
To show Corollary 1, we first obtain values of $\delta$ and $\epsilon$ such that the probability in Theorem 1 is $1-\varepsilon$. We use $\mathfrak{s}_{i},i=1,2,\ldots$ to denote different constants throughout this proof. For $a(n)=d_{1}/n$ with a sufficiently large $d_{1}$, we have $\beta_{n_{0}}(m)\leq 1/m$. This implies that
@@
Let $\varepsilon/(4d)=\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0})$, which gives us $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ for appropriate constants $\mathfrak{s}_{2}$ and $\mathfrak{s}_{3}$. This choice of $\delta$ gives us
$22d\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\varepsilon/$ 2 .
Using Theorem 1 from Korda and Prashanth (2015), we have $P(\|x_{n_{0}}-x^{*}\|\leq\epsilon)>1-\varepsilon/2$, where $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$. So, choosing $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ gives us
$P(\|x_{n_{0}}-x^{*}\|>\epsilon)\leq\varepsilon/2.$
Next we substitute the values $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon)$ and $\epsilon=(\mathfrak{s}_{4}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{5})n_{0}^{-1/2}$ in our bound from Theorem 1. Note that $\exp(-(1-\alpha)b_{n_{0}}(m-1))\epsilon\leq\epsilon$. Since $a(n)=d_{1}/n$, we have $a(n_{0})c_{2}=\mathfrak{s}_{6}n_{0}^{-1}$ and $a(n_{0})c_{1}\epsilon=(\mathfrak{s}_{7}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{8})n_{0}^{-3/2}$. Combining these terms, we get the following.
$e^{-(1-\alpha)b_{n_{0}}(m-1)}\epsilon+\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ $\leq\mathfrak{s}_{9}n_{0}^{-1/2}\log^{1/2}(1/\varepsilon)+\mathfrak{s}_{10}n_{0}^{-1/2}$
$=\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right).$
Hence we get that $\|x_{m}-x^{*}\|\leq\mathcal{O}\left(\left(\log^{1/2}(1/\varepsilon)+1\right)n_{0}^{-1/2}\right)$ holds for all $m\geq n_{0}$ with probability at least $1-\varepsilon$, completing the proof for Corollary 1.

### v3

@@ 1 Introduction
While there is a considerable body of work on concentration inequalities for reinforcement learning algorithms, these always aim for a moment bound or a high probability bound for sufficiently large time $n$. In contrast, we look at the complementary object, viz., concentration for all time after some $n_{0}\geq 0$. Moment bounds for TD(0) have been established, e.g., in Srikant and Ying (2019), Bhandari et al. (2018), and Chen et al. (2021). Other types of bounds (e.g., finite sample bounds with high probability results) have been established for multiple different settings in Prashanth et al. (2021) (consider uniform sampling from dataset), Patil et al. (2023) (consider projection and tail averaging), and Chen et al. (2025), Dalal et al. (2018a) (consider i.i.d. samples of state - action - next state triplets $(s,a,s^{\prime})$).
Going beyond TD(0), Bhandari et al. (2018), Chen et al. (2020, 2021), Wang et al. (2020), Dalal et al. (2018a, b), Li et al. (2020, 2023), Even-Dar and Mansour (2004), Prashanth et al. (2021), Qu and Wierman (2020), Sidford et al. (2018), Srikant and Ying (2019), Wainwright (2019a, b) present general high probability concentration bounds for various classes of algorithms. There is also a parallel body of work which seeks bound, either finite time or asymptotic, on the difference between the value function under the learned policy and the optimal value function (Jin et al. 2018, Yang et al. 2020, Yang and Wang 2019), i.e., the ‘regret’. Once again, this is distinct from our objective, which is to obtain a high probability bound valid for all time from some time on.
@@ 2 TD(0) - Background and Main Result
TD(0) is an algorithm for policy evaluation, i.e., for learning the performance of a fixed policy, and not for optimizing over policies (Tsitsiklis and Van Roy 1997). Hence a stationary policy is fixed a priori, giving us a time homogeneous uncontrolled Markov chain $\{Y_{n}\}$ over a finite state space $S$. The transition probabilities are given by $p(\cdot|\cdot)$, where the dependence on the policy is suppressed. Assume that the chain is aperiodic irreducible with the stationary distribution $\pi=[\pi(1),...,\pi(s)],s=|S|$. Let $D$ denote the $s\times s$ diagonal matrix whose $i$th diagonal entry is $\pi(i)$. Reward $r(i)$ is received when a transition from state $i$ takes place. The objective is to evaluate the long term discounted reward for each state given by the “value function”
@@
    and is assumed to be non-increasing, i.e., $a(n+1)\leq a(n)\ \forall\;n$. We also assume that $a(n)<1$ for all $n$. We further assume that $\frac{d_{1}}{n+1}\leq a(n)\leq d_{3}\left(\frac{1}{n+1}\right)^{d_{2}},\forall\ n$, where $d_{1}>0$ and $0<d_{2}\leq 1$. Larger values of $d_{1}$ and $d_{2}$ and smaller values of $d_{3}$ improve the main result presented below. The role this assumption plays in our bounds will become clear later.
@@
The following lemma shows that the function $\sum_{i}\pi(i)F(\cdot,i)$ is a contraction. Let $\langle x,x^{\prime}\rangle=x^{T}x^{\prime}$ and $\langle x,z\rangle_{D}=x^{T}Dz$. Note that the following lemma shows that $\sum_{i}\pi(i)F(\cdot,i)$ is contractive under the Euclidean norm with an explicit expression for the contraction factor.
@@
The term $P(\|x_{n_{0}}-x^{*}\|>\epsilon)$ captures the unavoidable contribution of the initial condition at $n_{0}$. This can be bounded by combining moment bounds (Srikant and Ying 2019, Bhandari et al. 2018, Chen et al. 2021) with Markov’s inequality.
@@
In Chen et al. (2025), an all-time bound is obtained which goes to zero as $m\uparrow\infty$. In our bound, the term $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$ remains constant as $m$ is increased. Here $\delta$ can be modified to $\delta(m)$ (similar to the treatment in Corollary 1 in Chandak et al. (2022)). But the term $a(n_{0})(c_{2}+c_{1}\epsilon)$ arises from our treatment of Markov noise using Poisson equation. Note that Chen et al. (2025) do not consider the case of Markov noise, but only consider the case of additive and multiplicative noise (i.i.d. samples of state-action-next state triplets). We leave incorporating their ideas into our approach to get a bound decaying with $m$ for Markov noise as a future work.
Remark 4.
For the special case of $a(n)=\frac{d_{1}}{n+1}$, we combine our result with Theorem 2.1 from Chen et al. (2021), a mean square error bound, to get the following corollary.
@@
Let $a(n)=d_{1}/(n+1)$ with sufficiently large $d_{1}$. Let $n_{0}$ be large enough to satisfy assumptions of Theorem 1 and (Chen et al. 2021, Theorem 2.1). Then with probability at least $1-\varepsilon_{1}-\varepsilon_{2}$, we have for all $m\geq n_{0}$ that
$\|x_{m}-x^{*}\|=\mathcal{O}\left(\frac{1}{\sqrt{n_{0}}}\log^{1/2}\left(\frac{1}{\varepsilon_{1}}\right)+\sqrt{\frac{\log(n_{0})}{n_{0}}}\frac{1}{\sqrt{\varepsilon_{2}}}\left(\frac{n_{0}}{m}+\frac{1}{n_{0}}\right)\right).$
The proof for this corollary has been presented at the end of Appendix B. The first term here corresponds to the term $\delta$ in Theorem 1. This term has a $\sqrt{1/n_{0}}$ decay rate and an exponentially small tail. The second term is the contribution of the initial condition at $n_{0}$. We have a polynomial tail in this case, but the dependence on $m$ and $n_{0}$ is stronger, as the term $(\sqrt{\log(n_{0})}/\sqrt{n_{0}})\times(n_{0}/m)$ decays with $m$ and the other term decays as $\log^{1/2}(n_{0})n_{0}^{-3/2}$.
@@ 3 Proof of the Main Result
Now we obtain a bound for $P\left(\bar{x}^{\prime}_{n,n-1}>\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}\right)$ by induction. Note that while $\bar{x}^{\prime}_{n-1,n-1}\leq\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$, $\bar{x}^{\prime}_{n,n-1}$ is not necessarily less than $\frac{a(n_{0})(c_{2}+c_{1}\epsilon)+\delta}{1-\alpha-a(n_{0})c_{1}}$. Hence when quantifying the above probability, we only have to look at $\|\bar{x}_{n,n-1}-\bar{z}_{n,n-1}\|$. Then,
@@
Then $A=\cap_{n=n_{0}}^{\infty}A_{n}$. Hence $P(A)=\lim_{n\uparrow\infty}P(A_{n})$. This completes the proof for Theorem 1. ∎
@@ 4 Conclusions
In conclusion, we note some future directions. The concept of relaxed martingale concentration inequalities can be used to obtain bounds of the similar flavor for algorithms which suffer from similar issues. These include TD($\lambda$) and SSP Q Learning. Alternatively, similar bounds can be obtained for variants of temporal difference learning (Chen et al. 2021). Another direction could be to improve the bounds in this paper to get an exponentially small tail for Markovian stochastic approximation.
@@ B.4 Proof of Corollary 1
To show Corollary 1, we first obtain values of $\delta$ and $\epsilon$ such that the probability in Theorem 1 is $1-\varepsilon_{1}-\varepsilon_{2}$. We use $\mathfrak{s}_{i},i=1,2,\ldots$ to denote different constants throughout this proof. For $a(n)=d_{1}/(n+1)$ with a sufficiently large $d_{1}$, we have $\beta_{n_{0}}(m)\leq 1/m$. This implies that
@@
Let $\varepsilon_{1}/(2d)=\mathfrak{s}_{1}\exp(-D\delta^{2}n_{0})$, which gives us $\delta=\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})$ for appropriate constants $\mathfrak{s}_{2}$ and $\mathfrak{s}_{3}$. This choice of $\delta$ gives us
$2d\sum_{m\geq n_{0}+1}\exp(-D\delta^{2}/\beta_{n_{0}}(m))\leq\varepsilon_{1}.$
Let $\epsilon=\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}/\sqrt{\varepsilon_{2}}$, which implies that
$P(\|x_{n_{0}}-x^{*}\|>\epsilon)=P\left(\|x_{n_{0}}-x^{*}\|^{2}>\frac{E[\|x_{n_{0}}-x^{*}\|^{2}]}{\varepsilon_{2}}\right)\leq\varepsilon_{2}.$
Here the last inequality follows from Markov’s inequality. Being linear contractive SA with an aperiodic irreducible Markov chain, our formulation satisfies the assumptions for Theorem 2.1 from Chen et al. (2021). To apply their result, we note the corresponding mapping between constants: the norm $\|\cdot\|_{c}$ is the Euclidean norm in our case, $h$ is $1$, $\varphi_{2}$ is $1-\alpha$ in our case and their $\alpha$ is $d_{1}$ in our case. For $d_{1}>1/(1-\alpha)$ and $n_{0}$ sufficiently large to satisfy the condition for Theorem 2.1 (2), we can use their result Theorem 2.1 (2)(b)(iii) to obtain the following mean square bound.
$E\left[\|x_{n_{0}}-x^{*}\|^{2}\right]$ $\leq\mathfrak{s}_{4}\left(\frac{1}{n_{0}+1}\right)^{(1-\alpha)d_{1}}+\mathfrak{s}_{5}\frac{\log(n_{0}+1)}{n_{0}+1}\leq\mathfrak{s}_{6}\frac{\log(n_{0}+1)}{n_{0}+1}.$
Substituting the values of $\delta$ and $\epsilon$ in our bound, we get with probability greater than $1-\varepsilon_{1}-\varepsilon_{2}$,
$\|x_{m}-x^{*}\|$ $\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\frac{\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}}{\sqrt{\varepsilon_{2}}}$
$\;\;+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\sqrt{E[\|x_{n_{0}}-x^{*}\|^{2}]}}{\sqrt{\varepsilon_{2}}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
$\leq e^{-(1-\alpha)b_{n_{0}}(m-1)}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}$
$+\mathfrak{s}_{7}\left(\frac{c_{2}d_{1}}{n_{0}+1}+\frac{c_{1}d_{1}}{n_{0}+1}\frac{\mathfrak{s}_{6}}{\sqrt{\varepsilon_{2}}}\sqrt{\frac{\log(n_{0}+1)}{n_{0}+1}}+\mathfrak{s}_{2}n_{0}^{-1/2}\log^{1/2}(\mathfrak{s}_{3}/\varepsilon_{1})\right)$
for all $m\geq n_{0}$. Now
$\exp(-(1-\alpha)b_{n_{0}}(m-1))$
$\leq\exp\left(-\sum_{i=n_{0}}^{m-1}(1-\alpha)a(i)\right)\leq\exp\left(-\sum_{i=n_{0}}^{m-1}\frac{(1-\alpha)d_{1}}{i+1}\right)$
[…]
$=\left(\frac{n_{0}+1}{m+1}\right)^{d_{1}(1-\alpha)}\leq\frac{n_{0}+1}{m+1}.$
Here the final inequality follows from the assumption that $(1-\alpha)d_{1}>1$. Hence we get that for sufficiently large $n_{0}$, the following holds with probability $1-\varepsilon_{1}-\varepsilon_{2}$ for all $m\geq n_{0}$.
$\|x_{m}-x^{*}\|=\mathcal{O}\left(\frac{1}{\sqrt{n_{0}}}\log^{1/2}\left(\frac{1}{\varepsilon_{1}}\right)+\sqrt{\frac{\log(n_{0})}{n_{0}}}\frac{1}{\sqrt{\varepsilon_{2}}}\left(\frac{n_{0}}{m}+\frac{1}{n_{0}}\right)\right).$